OpenStudy (kweb2017):

If f(x) = x^2 − 3x − 4 and g(x) = x^2 + x, what is (f + g)(x)?

OpenStudy (mathmale):

If f(x) = x^2 − 3x − 4 and g(x) = x^2 + x, what is (f + g)(x)? Line up f(x) and g(x), g(x) below f(x), and then add like terms together. f(x) = x^2 − 3x − 4 g(x) = x^2 + x ------------------

OpenStudy (kweb2017):

Thanks! It told me the answer was 2x^2-2x-4 2(x^2-x-2)

OpenStudy (mathmale):

Looks good. Be sure to label your result: (f+g)(x) =2( ... ).

OpenStudy (kweb2017):

Okay! Thank you!